OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8)

OpenStudy (saifoo.khan):

\[Slope = \frac{rise}{run} = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (saifoo.khan):

\[\frac{8-5}{6-4}\]